A uniform chain of length 2 m is kept on a table such that a length of 60 cm hangs freely from the edge of the table. The total mass of the chain is 4 kg. What is the work done in pulling the entire chain on the table? please explain each and every step thank u


Grade:11

3 Answers

suryakanth AskiitiansExpert-IITB
13 years ago

Dear aditya,

Given

M and L are total lenght and mass of the chain.
M = 4 kg
L = 2 m

And let length and mass of part of chain hanging down are l and m respectively

l = 0.6 m

considering mass per unit length =>

M/L = m/l

m = l / L * 4 = (0.6 / 2) * 4 = 1.2 kg

Tension in chain hanging down = mg = 11.76 N (g=9.8 m/s^2)

Force required = Tension
Displacement = 0.6 m (part of chain hanging comes up)
W= Force * Displacement = 11.76 * 0.6 = 7.056 J

Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

 

Askiitians Expert

Suryakanth –IITB

vikas askiitian expert
13 years ago

this q can be easily solved by considering motion of center of mass of hanging part of chain....

center of mass of hanging part lies in the middle ...

distance of center of mass from table is 30cm or 0.3m.....

workdone in pulling hanging part of chain  is equivalent to workdone in pulling center of mass of the hanging part to the table...

mass of comlete chain =4kg

 mass per unit length =4/2=2kg/m

 mass of 60 cm of chain=0.6*2=1.2kg

now workdone=mgH(com)

                    =1.2 * 10 *0.3

                    =3.6J

Rishi Sharma
3 years ago
Dear Student,
Please find below the solution to your problem.

Given
M and L are total lenght and mass of the chain.
M = 4 kg
L = 2 m
And let length and mass of part of chain hanging down are l and m respectively
l = 0.6 m
considering mass per unit length =>
M/L = m/l
m = l / L * 4 = (0.6 / 2) * 4 = 1.2 kg
Tension in chain hanging down = mg = 11.76 N (g=9.8 m/s^2)
Force required = Tension
Displacement = 0.6 m (part of chain hanging comes up)
W= Force * Displacement = 11.76 * 0.6 = 7.056 J

Thanks and Regards